OpenStudy (anonymous):

Jack knows the surface area of a cylinder and its radius. He wants to find the cylinder's height. He needs to rewrite the formula to find the cylinder's height (h) in terms of the cylinder’s surface area (A) and its radius (r). Which is the correct formula?

OpenStudy (instagrammodel):

work with me on this one.. A=2πrh+2πr² notes: what you do to one side of = sign, you must do to both sides. subtract the 2πr² A - 2πr² = 2πrh Now divide both sides by 2πr h = (A - 2πr²)/2πr This is a good answer. Or you could try to simplify more. we know that (a +b)/c = a/c + b/c h = (A - 2πr²)/2πr h = A/(2πr) -r

OpenStudy (anonymous):

a.h= r+ a/2Pir b.h=a/2Pir c.h=a/2Pir-r^2

OpenStudy (anonymous):

I honestly don't know im lost